two cars leave towns at the same time in oppoistite directions one travels 35 mi/h and the other 40 mi/h in how many hours will they be 15 miles apart?
Answer by mananth(16946)

Opposite directions away from each other
add up speeds
30+45=75 mph
t=d/r
15/75
=1/5 hour OR 12 minutes